### v2

# FastVAR: Linear Visual Autoregressive Modeling via Cached Token Pruning

Abstract
Visual Autoregressive (VAR) modeling has gained popularity for its shift towards next-scale prediction. However, existing VAR paradigms process the entire token map at each scale step, leading to the complexity and runtime scaling dramatically with image resolution. To address this challenge, we propose FastVAR, a post-training acceleration method for efficient resolution scaling with VARs. Our key finding is that the majority of latency arises from the large-scale step where most tokens have already converged. Leveraging this observation, we develop the cached token pruning strategy that only forwards pivotal tokens for scale-specific modeling while using cached tokens from previous scale steps to restore the pruned slots. This significantly reduces the number of forwarded tokens and improves the efficiency at larger resolutions. Experiments show the proposed FastVAR can further speedup FlashAttention-accelerated VAR by 2.7$\times$ with negligible performance drop of <1%. We further extend FastVAR to zero-shot generation of higher resolution images. In particular, FastVAR can generate one 2K image with 15GB memory footprints in 1.5s on a single NVIDIA 3090 GPU. Code is available at https://github.com/csguoh/FastVAR.
Figure 1: Our FastVAR can generate one 2K image using one NVIDIA 3090 GPU, while existing baseline fails due to out of memory.
1 Introduction
The next-token prediction of Autoregressive (AR) models [30, 44, 56, 32, 52] has demonstrated performance competitive with diffusion models for visual generation [44] as well as the potential for unified vision understanding and generation [56, 54, 52, 53]. However, this token-by-token paradigm suffers from numerous decoding steps. Recently, Visual Autoregressive (VAR) modeling [47] has shifted the paradigm to next-scale prediction, enabling image generation in fewer steps. Under this new paradigm, some works [21, 45] have developed VAR-based models for text-to-image generation and obtained promising results.
Despite their potential, existing VAR-based methods [47, 21, 45] face a critical challenge: the computational complexity and runtime latency scale dramatically with image resolution. Specifically, unlike next-token prediction which processes only one token per step, the next-scale prediction of VAR requires processing the entire token map at each decoding step. As a result, the number of tokens increases in $\mathcal{O}(n^{2})$ with the image resolution $n\times n$, and even leads to $\mathcal{O}(n^{4})$ complexity in the attention [50] layer. Empirically, as shown in  Fig. 2(a), even when FlashAttention [13] is enabled, VAR models still exhibit super-linear runtime latency. Consequently, this significant computational complexity prevents existing VAR-based models from scaling to higher resolutions, such as 2K.
Figure 2: FastVAR exhibits promising resolution scalability, and can achieve noticeable speedup with negligible performance drop.
In this work, we aim to address the resolution scaling challenge of VARs by pruning forwarded tokens. Through an in-depth analysis of the pre-trained VAR models, we identify the following key findings. 1) Large-scale steps are speed bottleneck but appear robustness. Runtime profiling in Fig. 2(a) finds that the last two large-scale steps account for 60% of the total runtime. Further investigation in Tab. 4 reveals that VAR is more resilient to token pruning at large-scale steps than at smaller scales. Thus, we direct our efforts to the large-scale steps. 2) High-frequency modeling matters at large-scale steps. Spectrum analysis in Fig. 3(b) reveals that the large-scale steps are optimized to model high-frequency tokens, such as texture details, and the remaining low-frequency tokens almost converge in these steps. Consequently, we can only forward the high-frequency tokens for pruning. 3) Tokens from different scales are related. Attention map analysis in Fig. 3(c) shows strong diagonal sparsity across scales, indicating that tokens attend not only to same-scale neighbors but also to those from the preceding scales. Thus, we cache tokens in previous steps to compensate for the pruned slots, preserving the 2D image lattice structure and maintaining information flow.
Based on the above observations, we propose FastVAR, a post-training acceleration recipe for efficient resolution scaling with VARs. At its core, FastVAR employs “cached token pruning”, which retains only pivotal tokens at large-scale steps to reduce computational overhead while using cached token maps from early-scale steps to compensate for information loss. To identify the pivotal tokens, we develop Pivotal Token Selection (PTS), a frequency-based scoring mechanism. Token importance is determined by filtering out the low-frequency component estimated via the direct current component, which enables efficient frequency-based token selection directly in the spatial domain. The selected Top-K pivotal tokens are then processed by the VAR model. To restore pruned tokens, we propose Cached Token Restoration (CTR), which first interpolates the token map from the cached scale and then reinstates the pruned tokens by indexing the interpolated token map to the pruned location. By integrating these strategies, we present FastVAR which enjoys the following benefits:
  • The proposed method is training-free and plug-and-play for various VAR-based backbones.
  • As shown in Fig. 2(b), FastVAR can be integrated with FlashAttention, with further 2.7$\times$ speedup and <1% performance drop.
  • In Fig. 1, FastVAR facilitates zero-shot scaling to larger-resolution and can produce one 2K image using only 15GB memory in 1.5s on a single NVIDIA 3090 GPU.
2 Related Work
Autoregressive Visual Generation. The previous autoregressive (AR) methods [61, 39, 26, 58] mostly adopt the next-token prediction paradigm, which treats each pixel as one token and generates pixels in a GPT or BERT style [6, 1, 34, 14, 4]. For instance, some pioneering works [48, 9] generate pixels in raster scan order with transformer models. Later, VQVAE [49] and VQGAN [15] propose to quantize image patches into discrete tokens to ease the training process. Benefiting from the scaling up [25, 23], recent works have shown promising results that the autoregressive models can achieve a more competitive performance than state-of-the-art Diffusion models [40, 38, 16, 8, 7]. Despite this progress, existing next-token generation pipelines struggle to efficiently synthesize high-resolution images due to the numerous decoding steps [10]. More recently, the Visual Autoregressive (VAR) modeling [47], which shifts to next-scale generation fashion to allow only a few forward steps, further advances the generation quality as well as the model efficiency. Building on top of VAR, existing methods have scaled up the model for production-level text-to-image generation. For example, HART [45] adopts the continuous diffusion module to compensate for the quantization error. Infinity [21] employs a bitwise tokenizer for an extremely large vector vocabulary. Despite the promising ability of VAR-based methods, they struggle to scale to larger resolutions with the increasing token numbers.
Efficient Visual Generation. To speed up diffusion models [24], many efforts have been made in recent years, including distillation [41, 37], quantization [43, 31], pruning [59, 2, 17, 62, 51], and caching [29, 33, 36, 35], which either reduce the total denoising steps or to reduce the cost of single forward. Specifically, DeepCache [36] proposes to reuse the intermediate features of low-resolution layers in the U-Net. ToMeSD [2] merges similar tokens into one token with subsequent unmerging for acceleration. Unfortunately, these methods are specifically designed for diffusion, which can achieve only sub-optimal performance or cannot be used in VAR. In the efforts to accelerate AR modeling [30, 44, 56, 32, 52], one prevalent solution is to reduce the forward step using parallel decoding strategies [22, 46, 42]. For instance, Speculative decoding [5, 27] utilizes a small draft model to generate candidate tokens, which are then verified in parallel by the larger model. However, the well-developed decoding strategy cannot be directly applied to the next-scale paradigm, whose each step involves multiple pixels [47]. As for fast VAR models, one related work is the CoDe [10], which uses model ensemble techniques to use a small model on the costly large resolution. However, CoDe highly depends on the availability of different-sized models. In this work, we attempt to speed up the VARs by pruning extra tokens, which is training-free and is generic for multiple VAR methods.
Figure 3: (a) We resize different $\tilde{r}_{k}$ to the same size for better presentation. (b) Each curve represents the frequency characteristics of a certain-scale token map. (c) The row represents keys of scale steps $\{1,2,\cdots,k\}$, and the column is queries of the $k$-th scale step.
3 Method
3.1 Preliminary
The Visual Autoregressive (VAR) modeling [47] redefines the traditional AR by shifting from a “next-token prediction” to a “next-scale prediction”, where each autoregressive unit is a token map of varying scales instead of a single token. For a given image feature map, VAR first quantizes it into $K$ multi-scale token maps =$\mathcal{R}=\{r_{1},r_{2},\cdots,r_{K}\}$ with progressively larger resolutions using pre-definced scale schdual {$(h_{1},w_{1}),(h_{2},w_{2}),\cdots,(h_{K},w_{K})$}. Then the $\mathcal{R}$ is used to train a Transformer [50] to learn the joint distribution using the causal formulation:
$p(r_{1},r_{2},\cdots,r_{K})=\prod_{k=1}^{K}p(r_{k}|r_{1},r_{2},\cdots,r_{k-1}),$ (1)
where the start token $r_{1}$ is the encoded text embedding from the LLMs [57, 11] under given user prompts. During training, existing methods [45, 21, 47] usually employ a residual strategy to reduce learning difficulty. In detail, the model outputs at the $k$-th scale step $f_{k}$ is the residual of the intermediate prediction $\tilde{r}_{k}$:
$\tilde{r}_{k}=\mathrm{interpolate}(\tilde{r}_{k-1},(h_{k},w_{k}))+f_{k},$ (2)
where the “$\mathrm{interpolate}(\cdot,(h_{k},w_{k}))$” denotes upsample a token map to the size of $(h_{k},w_{k})$. Unfolding  Eq. 2 derives the following cumulative form:
$\tilde{r}_{k}=\sum_{i=1}^{k}\mathrm{interpolate}(f_{i},(h_{k},w_{k})).$ (3)
In the last scale step, the predicted token maps $\tilde{r}_{K}$ is used to generate the final image. During inference, since the outputs at different scales are generated autoregressively, the KV-Cache can be adopted to avoid re-computing of previous steps. Although the VAR paradigm can generate images within only $K$ scale steps, the Transformer needs to process all $h_{k}\times w_{k}$ tokens at the $k$-th step, hindering scaling to higher-resolution image synthesis.
Figure 4: (a) We keep the Structure Construction Stage as standard VAR and additionally store token maps at the $(K-N)$-th step as the cache. Then the Texture Filling Stage applies FastVAR to process only pivotal tokens for fast inference. (b) For Attention or FFN layers using FastVAR, we first employ the Pivotal Token Selection (PTS) for token reduction. After the model forward, we develop the Cached Token Restoration (CTR) to restore the original token numbers using the caches from the previous scale step.
3.2 Motivation
In this work, we aim to approach linear VAR models to advance high-resolution image generation. To this end, we investigate the scale-wise behavior of the pre-trained VAR model [21]. We examine from three angles: the latency profiling, the specific role of each step, and the cross-scale token dependency. Conclusions are summarized as follows.
Observation 1: Large-scale steps are speed bottleneck but appear robustness. As shown in the runtime curve in Fig. 2(a), even the FlashAttention is enabled, the inference latency of the VAR exhibits a super-linear complexity with resolution steps. This is because the Transformer needs to take all the tokens at once. As a result, the tokens from the large-scale steps make up the majority of the total token length. For example, the last two steps even occupy 60% of the total time. On the other hand, the experiments in Sec. 4.3 suggest the large-scale steps are more robust to token drop than small-scale counterparts under the same pruning ratio. Therefore, it is accessible to prune some “unimportant tokens” at the large scale step to speed up token processing. However, challenges still remain about which tokens should be pruned and how to compensate for the information loss from these pruned tokens.
Observation 2: High-frequency modeling matters at large-scale steps. To find which token should be pruned, we visualize the intermediate predictions $\tilde{r}_{k}$ in Fig. 3(a). Interestingly, we find that the generation process of the pre-trained VAR model can be decomposed into two stages. The Structure Construction Stage, which includes small-scale steps, is trained to generate the outline of a subject. After that, at the Texture Filling Stage, which is composed of large-scale steps, is responsible for adding details based on the previous sketch. Given that the outline is low-frequency while the texture is high-frequency, we therefore infer that the VAR mainly models the high-frequency texture in the large-scale stage while keeping the low-frequency structure almost intact. For further verification, we show in Fig. 3(b) that the low-frequency component almost converges in the large-scale steps, while the high-frequency modeling stays optimized with noticeable variations. Therefore, we can prune the redundant low-frequency tokens and forward only the pivotal high-frequency tokens at the large-scale steps for token number reduction.
Observation 3: Tokens from different scales are related. After reducing the input tokens for fast forward, we still need to restore the output tokens to the original numbers in order to maintain the 2D image structure. To this end, we analyze the cross-scale attention map of the unpruned VAR model, where the query at the current scale interacts with KV caches from all previous scales. As depicted in Fig. 3(c), we observe that the attention map not only exhibits a high diagonal score at self-scale, but also a high diagonal score at cross-scale. This behavior indicates that a token in the current scale not only attends to its same-scale locals but also exhibits strong correlation to tokens at the neighboring positions in previous scales. This diagonal attention sparsity thus facilitates us to use only a few tokens to estimate the original output of the pruned slots, and avoids weighting multiple tokens across the token map. Specifically, we can approximate the outputs of pruned tokens by indexing in the previous-scale token maps based on the pruning location. In this way, the information loss can be compensated thanks to this high cross-scale similarity.
3.3 Efficient Resolution Scaling for VARs
We instantiate the above idea and present FastVAR to approximate linear VAR models. The core is the cached token pruning, which prunes tokens at large scale steps while using cache from early scales to maintain information flow.
Overview. As shown in Fig. 4(a), denote the scale step set of Structure Construction Stage as $\mathcal{S}=\{1,2,\cdots,K-N\}$, the step set of Texture Filling Stage as $\mathcal{T}=\{K-N+1,\cdots,K\}$. As stated in Sec. 3.2, the speedup of pruning token maps in $\mathcal{S}$ is insignificant while producing a noticeable performance drop, we therefore keep $\mathcal{S}$ intact as the standard VAR. In addition, we save the per-layer outputs at the last scale in $\mathcal{S}$, i.e., the $(K-N)$-th step, as the cache for subsequent token restoration. For the set $\mathcal{T}$, we apply FastVAR to prune tokens for fast-forward, followed by the token number restoration. More details are given below.
Pivotal Token Selection. We first introduce the Pivotal Token Selection (PTS) to reduce the number of tokens. As observed in Fig. 3(b), the low-frequency tokens almost converge in the later scale steps. Therefore, we can feed only the high-frequency tokens and prune the remaining. However, the challenge lies in that common frequency operators, such as FFT, work in the frequency domain, making it difficult to recognize the frequency characteristics of certain tokens in the original token map. To this end, our proposed PTS adopts an approximation solution. Specifically, denote $x_{k}$ as the input of one layer, we estimate its low-frequency component $\bar{x}_{k}$ as the direct current component, which can be easily obtained via the spatially global average pooling:
$\bar{x}_{k}=\mathrm{global\_avg\_pool}(x_{k}).$ (4)
After that, the high-frequency component is computed as the difference between the $x_{k}$ and $\bar{x}_{k}$. The pivotal score $s_{k}$ is defined as the L2 norm of the high-frequency maps:
$s_{k}=||x_{k}-\bar{x}_{k}||_{2}.$ (5)
Subsequently, we obtain the index set $\mathcal{I}$ for pivotal token selection by keeping TopK scores in $s_{k}$. Finally, we perform token pruning on $x_{k}$ through indexing with $\mathcal{I}$ to allow the fast model forward. Notably, since the PTS reduces the number of input tokens in one transformer layer, as an additional benefit, the KV-cache is also reduced accordingly, thus optimizing the GPU footprints as well as the cross-scale attention for the subsequent scale steps.
Cached Token Restoration. The proposed PTS can effectively reduce the number of tokens. However, the visual generation tasks need to restore the original token numbers to regain the structural 2D image. For this reason, we propose Cached Token Restoration (CTR). The main rationale stems from the diagonal attention sparsity that the token at the $k$-th scale not only attends to itself, but also exhibits strong correlation with the corresponding slot from previous $\{1,2,,\cdots,k-1\}$ steps. Therefore, our proposed CTR approximates the layer outputs in the pruned locations using the counterparts from the cached step. Formally, given the cached token map $y_{K-N}$ with size $(h_{K-N},w_{K-N})$, which is the output of the corresponding layer at step $K-N$, we first upsample it to the same size of the current scale:
$y^{cache}_{k}=\mathrm{interpolate}(y_{K-N},(h_{k},w_{k})).$ (6)
Then the value from $y^{cache}_{k}$ is scattered into the pruned slots of $y_{k}$ using the index set $\mathcal{I}$, to generate the restored layer output $y^{\prime}_{k}$, which shares the same token numbers as the $x_{k}$. We also provide an explanation for the choice of $K-N$ as the caching step, see Suppl. for more details.
Implementation of FastVAR. Thanks to the token number invariance of the PTS+CTR, as shown in Fig. 4(b), we apply this operator pair for each Attention&FFN layer at the large-scale steps. Furthermore, we empirically find in Sec. 4.3 that the larger scale steps exhibit stronger robustness to pruning than that of early scale steps. Therefore, we develop a progressive pruning ratio schedule, in which we assign larger pruning ratios to larger scale steps in the set $\mathcal{T}$. We summarize the complete algorithm of FastVAR in the Suppl.. At last, it is worth noting that our FastVAR does not access the attention map, and we show in Sec. 4.4 that the proposed FastVAR can be used in combination with other acceleration techniques like FlashAttention [13, 12] for even faster VAR generation.
4 Experiments
4.1 Experimental Setup
Models and Evaluations. We apply our FastVAR on two VAR-based text-to-image models, namely HART [45] and Infinity [21]. Both models can generate images of up to $1024\times 1024$ resolution. For a fair comparison, we keep all the hyperparameters the same as their default settings. For evaluation metrics, we compare both in terms of generation quality and inference efficiency. For generation quality, we use two popular benchmarks, i.e., the GenEval [20] and MJHQ30K [28] to validate the high-level semantic consistency and the perceptual quality, respectively. For efficiency evaluation, we adopt metrics including running time, throughputs, speedup ratio, and GPU memory costs.
Implementation Details. For the number of pruned scale steps $N$, we set $N=4$ for Infinity and $N=2$ for HART, i.e., only token maps from the last 4 or 2 scale steps are applied with the proposed FastVAR, with other steps kept as standard VAR. For the progressive pruning ratio schedule, we set it to {40%, 50%, 100%, 100%} for Infinity, and {50%, 75%} for HART. The 100% pruning ratio indicates all tokens are dropped, i.e., we skip the corresponding steps and interpolate the intermediate outputs to the target resolution as the final outputs. Note that this extreme ratio depends on the pruned backbones, and we provide further discussion in the Suppl.. Unless specified, the efficiency of the unpruned baselines are already accelerated by FlashAttention [13], and we apply the proposed FastVAR on top of it. Following the setup of existing methods [47, 44, 10], the inference speeds for all methods are measured without including the VAE time cost since this is a shared cost for all methods. All experiments are conducted on one single NVIDIA RTX 3090 GPU with 24GB memory.
4.2 Main Results
Table 1: Quantitative comparison on efficiency and quality on 1024$\times$1024 GenEval benchmarks. The marks $\bullet$, $\circ$, and $\diamond$ denote the Diffusion, AR, and VAR-based methods, respectively. Note that the efficiency of HART and Infinity baselines are tested under FlashAttention.
Methods Inference Efficiency Generation Quality
[…]
$\diamond$+ FastVAR 14 1.5$\times$ 0.63s 1.59it/s 0.7B 0.57 0.16 0.24 0.51
[…]
$\diamond$+ FastVAR 13 2.7$\times$ 0.95s 1.05it/s 2.0B 0.81 0.45 0.52 0.72
Figure 5: Qualitative comparison between the original baselines and our proposed FastVAR on 1024$\times$1024 image generation. Our FastVAR achieves significant speedups, e.g., 2.7$\times$ on Infinity [21], while keeping high-quality results similar to the original model.
Comparison on GenEval. We first evaluate the quality-efficiency trade-off on $1024\times 1024$ text-to-image generation using the GenEval benchmark [20]. We compare our FastVAR against various state-of-the-art methods, including Diffusion, AR, and VAR models. The results are shown in Tab. 1. Compared with traditional AR models, our Infinity+FastVAR can achieve even 39.7$\times$ speedup than LlamaGen [44], while achieving 125% performance boosts on the GenEval. Moreover, our FastVAR achieves acceleration almost without performance loss, e.g., 1.5$\times$ speedup on the HART with the same GenEval score as the unpruned baseline. Our approach also reduces the inference GPU memory. For example, FastVAR achieves 22.2% reduction to 14.7GB on top of FlashAttention of 18.9GB, to produce a 1024$\times$1024 image in 0.95s, facilitating generation on consumer-level GPUs. The above results demonstrate the generality and effectiveness of our method.
Comparison on MJHQ30K. In Tab. 2, we further validate the perceptual quality on the MJHQ30K [28] benchmark. It can be seen that our FastVAR achieves a reasonable performance while maintaining a high speedup ratio. For instance, on the well-known challenging “people” category, our HART+FastVAR even achieves a FID reduction of 2.42 with 1.5$\times$ acceleration. In other categories, our method also maintains good performance with significant acceleration. We also give a qualitative visualization in Fig. 5, and one can see that the generated images from FastVAR maintain exceptionally high quality and accurate semantic information, indicating the effectiveness of the proposed method.
Figure 6: Quantitative and qualitative ablation experiments on different pruning ratios. We set varying pruning ratios on the last scale step of the HART model while keeping the other scale steps as the original setup.
Table 2: Quantitative comparisons of FID and CLIP score with different generation categories on the MJHQ30K dataset.
Methods Speedup landscape people food
[…]
SDXL [38] - 30.78 26.35 35.56 28.01 35.26 27.98
[…]
+ FastVAR 1.5$\times$ 22.52 26.51 28.19 28.34 30.97 28.25
[…]
+ FastVAR 2.7$\times$ 24.68 26.62 30.55 28.28 32.54 27.08
Difference from ToMe. Token Merging (ToMe) [3], which merges multiple tokens into one token for efficient forward followed by token unmerging to restore the 2D shape, appears applicable to accelerate VAR backbones due to its token number invariance. In Tab. 3, we set different FastVAR and ToMe variants using varying pruning ratios. One can see that ToMe fails to achieve high speedup. For example, even the 1.36$\times$ speedup can lead to noticeable FID degradation since it is difficult to compress the whole token map into limited tokens. In contrast, our FastVAR can achieve 1.7$\times$ higher speedup with better FID performance.
Table 3: Quantitative comparison results with ToMe [2] under different speedup settings.
[…]
HART [45] 1.00$\times$ 0.95s 1.05it/s 30.61 28.47 0.51
[…]
+ ToMe [2] 1.36$\times$ 0.70s 1.43it/s 35.22 27.99 0.46
[…]
+ FastVAR 1.70$\times$ 0.56s 1.79it/s 28.97 28.24 0.50
Table 4: Ablation experiments on the scale-wise sensitivity. We focus on the scale range [16,21,27,36,48,64] with every two consecutive scales as a pruning group. The FastVAR is applied to the selected group while keeping the others unpruned.
[…]
baseline 0 1.00$\times$ 0.95s 30.61 28.47 0.51
[…]
$[27,36]$ 50% 1.09$\times$ 0.87s 27.92 28.49 0.51
$[27,36]$ 75% 1.14$\times$ 0.83s 29.38 28.56 0.50
$[48,64]$ 50% 1.30$\times$ 0.73s 27.13 28.43 0.51
$[48,64]$ 75% 1.64$\times$ 0.58s 28.56 28.36 0.50
4.3 Ablation Studies
Different Pruning Ratios. The pruning ratio plays an important role in balancing between efficiency and performance. Here, we conduct ablation experiments to investigate the impact of different pruning ratios. Both the quantitative and qualitative results are given in Fig. 6. Intuitively, increasing the pruning ratio leads to a stable runtime reduction since the model only needs to process fewer tokens. However, an over-large pruning ratio can also bring performance degradation due to the information loss of pivotal high-frequency tokens. Furthermore, the quantitative visualization suggests that increasing the pruning ratio makes certain textures and details discontinuous since the cache from the previous scale is not always an optimal approximation of the pruned tokens. Therefore, a moderate pruning ratio, e.g., 40%-75%, is adopted for HART+FastVAR to strike a sweet spot between performance and efficiency.
Scale-wise Sensitivity. As discussed in Sec. 3.2, we only focus on the token pruning of the last few scale steps. Here, we conduct ablation experiments in Tab. 4 to justify the rationality. It can be seen that the acceleration from pruning at an early scale steps is very limited due to the small token map size. For example, even reducing 50% tokens at the 48 and 64 scale steps is significantly faster than that of 75% at 16 and 21 scale steps. In addition, pruning at early scales also leads to a noticeable performance drop on both perceptual quality and semantic consistency. For instance, pruning at 48 and 64 scale steps brings 5.71 lower FID compared to that of 16 and 21 scale steps. This is because token reduction at small-scale steps disrupts the structure construction stage, and the inaccurate subject structure can exacerbate errors in subsequent scale steps.
Figure 7: Visualization of pruned tokens with different ratios.
Figure 8: We extend the scale schedule of pre-trained VAR model [21] for zero-shot larger resolution generation. We chose a modest extended scale since a larger one is observed quality drop due to the resolution gap. The images are re-scaled for better presentation.
Visualization of Pruned Slots. In Fig. 7, we visualize the selected TopK pivotal tokens under different pruning ratios. For a given pruning ratio, e.g., 50%, our FastVAR prioritizes the retention of high-frequency edges or texture slots, and removes tokens from flat regions that have already converged at early scale steps. Furthermore, the importance order within the selected tokens is also meaningful. For example, in the first row, when increasing the pruning ratio from 50% to 90%, the human eyes, hair and mouth regions, which are the most detailed, are consistently preserved while the low-frequency cheek regions are gradually pruned. Given that our FastVAR is gradient-free, this quantitative result confirms that the proposed frequency-based pivotal token selection is a reasonable heuristic.
4.4 Discussion
In Conjunction with FlashAttention. Existing token pruning methods [60, 19, 18, 55, 62] often depend on the attention map for token importance, leading to difficulties in combination with FlashAttention [13] in which the attention map is inaccessible. In comparison, our FastVAR can be seamlessly integrated with FlashAttention to facilitate more efficient inference. In Tab. 5, we show that using the FastVAR-only can obtain even 2.1$\times$ speedup against FlashAttention only, with 20.4% GPU memory reduction. After combining both, the resulting version leads to 2.7$\times$ speedup than FlashAttention only. This experiment validates that our FastVAR can be orthogonally combined with other methods for further acceleration.
Towards Larger Resolution. As presented in Sec. 3.2, the original VAR struggles to scale to higher-resolution due to the increasing token numbers. Our FastVAR can facilitate efficient resolution scaling, offering the potential to generate images at a larger resolution. To this end, we apply our FastVAR to the Infinity [21] for zero-shot larger resolution synthesis by appending additional steps. The results are shown in Fig. 8. It is noteworthy that even the FlashAttention accelerated baseline is out of memory on a 24GB NVIDIA 3090 GPU. In contrast, our FastVAR costs only 15GB memory and 1.3s runtime to generate a 1344$\times$1344 image. Moreover, the generated images are of high quality, suggesting that our FastVAR can facilitate production-level image generation on consumer-level GPUs.
Table 5: Efficiency comparison with FlashAttn on 1024$\times$1024 image generation.
[…]
SlowAttn [50] - - - OOM
FlashAttn only [13] 1.0$\times$ 2.61s 0.38it/s 16.1GB
SlowAttn+FastVAR 2.1$\times$ 1.25s 0.80it/s 12.8GB
[…]
5 Conclusion
This work presents FastVAR to advance high-resolution image synthesis with VARs by addressing the resolution scaling challenge. We reveal that the main latency bottleneck is the large-scale steps at which the low-frequency tokens have almost converged. To this end, we develop the cached token pruning technique to allow only process pivotal high-frequency tokens and use caches from the previous scale steps to restore the original token numbers. Thanks to the reduced forwarded tokens, our FastVAR approximates linear complexity, further enabling generation at a larger resolution such as 2K. Extensive experiments validate our FastVAR as an effective and generic solution for efficient resolution scaling of VAR models.